solve the triangle
a=5, b=12, c=10
What degrees is A, B and C?

Answer by stanbon(75887)

solve the triangle
a=5, b=12, c=10
What degrees is A, B and C?
-------
Law of Cosines:
a^2 = b^2+c^2-2bcCos(A)
----
Cos(A) = (b^2 + c^2 - a^2)/2bc
----
Cos(A) = (12^2+10^2-5^2)/(2*12*10) = 0.9125
===
A = cos^-1(0.9125) = 24.15 degrees
---------
Now use Law of Sines to find B:
sin(B)/b = sin(A)/a
---
sin(B) = (12/5)sin(24.15) = 0.9819
B = 79.08 degrees
-------------------------------
Then, C = 180-(24.15+79.08) = 76.77 degrees
============================
Cheers,
Stan H.
============================